# Exploring Channel Distinguishability in Local Neighborhoods of the Model Space in Quantum Neural Networks

### v1

@@ Abstract
With the increasing interest in Quantum Machine Learning, Quantum Neural Networks (QNNs) have emerged and gained significant attention. These models have, however, been shown to be notoriously difficult to train, which we hypothesize is partially due to the architectures, called ansatzes, that are hardly studied at this point. Therefore, in this paper, we take a step back and analyze ansatzes. We initially consider their expressivity, i.e., the space of operations they are able to express, and show that the closeness to being a 2-design, the primarily used measure, fails at capturing this property. Hence, we look for alternative ways to characterize ansatzes by considering the local neighborhood of the model space, in particular, analyzing model distinguishability upon small perturbation of parameters. We derive an upper bound on their distinguishability, showcasing that QNNs with few parameters are hardly discriminable upon update. Our numerical experiments support our bounds and further indicate that there is a significant degree of variability, which stresses the need for warm-starting or clever initialization. Altogether, our work provides an ansatz-centric perspective on training dynamics and difficulties in QNNs, ultimately suggesting that iterative training of small quantum models may not be effective, which contrasts their initial motivation.
@@ 1 Introduction
In light of these issues, we hypothesize that the architectures, called ansatzes, are part of the problem. QNN ansatzes consist of fixed and trainable gates. However, due to hardware limitations, mainly so-called Hardware Efficient Ansatzes (HEAs) (Kandala et al., 2017) are used, that are built primarily considering hardware constraints. Therefore, we hypothesize that executing the circuit does not lead to a meaningful feature representation.
@@
In particular, in classical ML, a vanilla CNN is ultimately designed to consider local features in lower layers, before considering global features in deeper layers, independently of the dataset. Transformers (Vaswani, 2017) work similarly, however, the trainable one-qubit and fixed two-qubit gates that are used in QNNs do not allow drawing intuitive conclusions about how features are extracted.
Therefore, the goal of this paper is to analyze and characterize QNN architectures. We aim to take a first step at exploring ansatzes and the model space, initially by considering their expressivity. Using tools from Quantum Information Theory and extensive numerics, we later analyze how the quantum circuit, i.e., the applied operation, changes upon parameter update, thus, providing insights into the model space. Our contributions are as follows.
  • Based on literature, we argue that closeness to a 2-design, a widely used measure for expressivity, does not adequately represent the possible set of operations in the unitary group.
@@
  • We provide numerical evidence that each ansatz is hardly distinguishable from itself upon parameter update during training, even in early stages, when larger updates occur.
  • The observed behavior has remarkable similarities to Barren Plateaus, thus, insinuating that the architectures, independently of the data or loss used, may be flawed and play a fundamental role in the trainability problems observed today.
The paper is structured as follows. In Section 2, we provide background information on QNNs, existing expressivity measures and related work [^1] For those unfamiliar with quantum computing, we refer to Appendix A.1 for a brief overview.. In Section 3, we discuss the closeness to being a 2-design as a measure of expressivity and Section 4 introduces a measure of distinguishability of quantum channels. We elaborate on our experimental setup in Section 5, and present our results in Section 6. Finally, Section 7 discusses implications of our results and Section 8 concludes the paper.
@@ 2.1 Quantum Neural Networks
The small system sizes and the noise in today’s Noisy Intermediate-Scale Quantum (NISQ) technology (Preskill, 2018) prevents execution of large-scale circuits. Consequently, so-called Variational Quantum Circuits (VQC) are applied, which are hybrid models composed of classical and quantum parts. For ML, they are now being referred to as QNNs due to similar optimization than classical ones, despite being rather related to kernel methods (Schuld, 2021).
Simply put, a QNN consists of (1) an input state |$\ket{\psi_{\bm{X}}}$, (2) an ansatz $U(\bm{\vartheta})$, (3) a loss function, and (4) an optimizer. The features are encoded onto the quantum state by means of a unitary feature encoding method $V(\bm{X})$, and the initial state is obtained as |$\ket{\psi_{\bm{X}}}=V(\bm{X})\ket{0}^{\otimes n}$, with $n$ as the number of qubits. HEAs consists of $L$ layers of trainable one-qubit rotation gates $V_{i}(\vartheta_{i})$ and fixed two-qubit entangling gates $W_{i}$, with $i\in[1,L]$, expressed as Equation 1. A measurement is taken in the end, which is done through a Hamiltonian observable $\hat{O}$, and the result is expressed as the expectation value of the ansatz with respect to |$\ket{\psi_{\bm{X}}}$ and $\hat{O}$, as shown in Equation 2.
@@
2.2 Expressivity of QNNs
The expressivity of a QNN is mainly defined as its ability to represent a wide range of functions (Tao et al., 2024). While several measures for expressivity have been proposed (Du et al., 2022; Schuld et al., 2021; Abbas et al., 2021), the most common method is to evaluate the similarity between the probability distribution of the unitaries generated by the ansatz and the Haar measure [^2] We refer to Appendix A.3 for a short introduction to unitary groups and the Haar measure., namely the uniform distribution over all unitaries $U(d)$ (Cerezo et al., 2021a). When the probability distribution of the unitaries follows the Haar measure up to the $t$-th moment, an ansatz is considered a t-design. Constructing such an ansatz requires exponential time (Harrow and Low, 2009), however, approximate t-designs can be built efficiently (Dankert et al., 2009).
@@ 2.3 Related Work
BPs have been proven for deep (McClean et al., 2018) and shallow circuits with global measurement (Cerezo et al., 2021b), expressivity (Holmes et al., 2022; McClean et al., 2018), noise (Wang et al., 2021), entanglement (Ortiz Marrero et al., 2021) and it has been linked to cost concentration and narrow gorges (Arrasmith et al., 2022). Further, even shallow models without BPs have only a small fraction of local minima close to the global minimum (Anschuetz and Kiani, 2022) [^3] For further information, we refer to  Larocca et al. (2024), who published an excellent review on the topic..
The problem has been linked to the curse of dimensionality in Cerezo et al. (2024), therefore, limiting the accessible Hilbert space was proposed. However, they showed that such models can be simulated classically, given an initial data acquisition phase on a quantum computer. This was extended to a trade-off between trainability and dequantizability, i.e., the potential to match the performance using classical models, in Gil-Fuster et al. (2024).
Besides that, parameter initialization strategies have been proposed, with some even providing better bounds on gradient magnitudes (Rad et al., 2022; Wang et al., 2023; Kulshrestha and Safro, 2022; Grant et al., 2019), however, the results of Herbst et al. (2024) imply that it is not the type of statistical distribution, but rather the used parameter ranges that could lead to better starting points.
Further, expressivity is commonly discussed to prove a quantum advantage with respect to classical models. The closeness to being a 2-design, as discussed in Section 2.2, is the most widely employed measure, however, others have been proposed. In particular, Equation 3 is directly related to the Frame Potential (Gross et al., 2007; Roberts and Yoshida, 2017; Nakaji and Yamamoto, 2021). Another approach is to analyze the entangling capability as in Sim et al. (2019), i.e., looking at the entanglement of states produced by random unitaries in the ensemble. Other measures are the Covering Number from  Du et al. (2022), analyzing Fourier coefficients (Schuld et al., 2021; Caro et al., 2021), or the Effective Dimension from Abbas et al. (2021). Another approach is to look at unitary t-designs, i.e., only consider the operations without the input state [^4] Unitary designs and the difference between unitary and state t-designs are discussed in Appendix A.3.. We employ the closeness to being a 2-design in our study, as it is frequently used, due to its intuitive interpretation. While there is a lot of research on measuring expressivity of a whole model, i.e., including data, measurement and loss, there is very little research on the architectures themselves.
@@
Changes in quantum channel or loss landscape upon perturbing parameters has not yet been studied for QNNs. However, perturbation analysis and sensitivity analysis have emerged as tools in classical ML. Examples are perturbing the input for explainability (Ivanovs et al., 2021; Pizarroso et al., 2022; Fel et al., 2023), or checking for stability (Testa et al., 2024). Similarly, weights can be perturbed for trainability (Wen et al., 2018), or to increase robustness and generalization (Wu et al., 2020; He et al., 2019; Dumford and Scheirer, 2020). To the best of our knowledge, there is no research on perturbing parameters to check how the underlying function changes, as, due to the stochasticity of quantum computing, this research question seems to be much more relevant in the quantum realm.
@@
While 2-designs have turned out to be sufficient for many quantum information processing protocols (Holmes et al., 2022; Harrow and Low, 2009), in this section, we summarize and connect results from the literature to argue that closeness to a 2-design is an inadequate measure for model expressivity (i.e., the ability to sufficiently approximate the unitary group to a high degree).
@@
Quantum state t-designs reduce integrals of polynomials over all quantum states to averages over a discrete set (cf. Equation 20). These are probability distributions over pure quantum states that replicate properties of the uniform (Haar) measure on the quantum states up to the $t$-th moment.
We measure expressivity of 2-design based architectures by comparing (generalized) Welch bound (Scott, 2008) values for state 2-designs against its t-design counterparts, with $t>2$. We state the inequality for a state t-design based architecture:
State t-design Welch bound: The choice of the polynomial function for state t-designs is presented in Appendix A.4. We state the following inequality (cf. Appendix B.2 for the derivation):
$\sum_{1\leq j,k\leq n}|\langle\bm{\psi}_{j}|\bm{\psi}_{k}\rangle|^{2t}$ $\geq\frac{n^{2}}{\binom{d+t-1}{t}}$ (4)
$=\frac{n^{2}}{\frac{(d+t-1)(d+t-2).....d}{t!}}\ \ \forall t\in\mathbb{N}.$
It is clear that for an increasing $t$, the term in the denominator $\frac{(d+t-1)(d+t-2).....d}{t!}$ increases rapidly (as compared to Equation 26 for 2-designs), yielding a vanishing overlap between the state vector summands. This makes the inequality in Equation 4 more stringent for increasing $t$ values. This indicates that higher-order designs ($t>2$) approximating the Haar measure up to the $t$-th order impose stronger constraints, requiring greater separation between a sequence of state vectors $\{\bm{\psi}_{1},\bm{\psi}_{2},....,\bm{\psi}_{d}\}$ to maintain equality in the bounds. Intuitively, states forming $t$-designs for $t>2$ are more “sparsely distributed” compared to their 2-design counterparts. This means that 2-design models have far fewer degrees of freedom in building complex quantum circuits. On the contrary, t-designs ($t>2$) can efficiently construct higher-order representations capturing more complex interactions between quantum states. Thus, evaluating Welch bounds for $2$-designs corroborates its inadequacy for constituting an expressive architecture.
Currently, there is no consensus in the community about a good measure of expressivity; indeed, it seems to be an open issue to define what even constitutes a highly expressive ansatz. Further, considering that the unitaries grow exponentially in system size, concepts in the Quantum Information Theory literature are often not efficiently computable, adding another layer of complexity to defining a useful measure. Additionally, there is evidence that trainability issues go hand-in-hand with architectures that span large parts of the $SU(N)$, i.e., are expressive with respect to the set of unitaries. Hence, it is questionable whether finding an alternative measure would help in practice in finding trainable ansatzes. Therefore, we adopt a different methodology and study how the QNN changes upon perturbation of parameters, which allows analyzing training dynamics and tracking the extent to which the QNN changes during training.
@@ 4 Channel Sensitivity
The diamond norm of a superoperator $T$ is defined as (Harrow and Low, 2009; Kitaev et al., 2002).
@@
Operationally, the diamond norm measures the maximum distinguishability between the output states of the two maps under any input state, i.e., for any input state we will be less or equally able to distinguish the output. A value of 00, means that they are indistinguishable, whereas a value of $2$ means they are perfectly distinguishable. Computing the diamond norm is non-trivial, however, can be reduced to a semi-definite program (Watrous, 2009, 2013). We apply the diamond norm to evaluate the distinguishability of a parametrized unitary $U(\bm{\vartheta})$, representing a quantum channel, from the unitary $U(\bm{\vartheta}+\bm{\delta})$, where $\bm{\delta}$ is a perturbation of $\bm{\vartheta}$, as follows.

### v2

@@ Abstract
With the increasing interest in Quantum Machine Learning, Quantum Neural Networks (QNNs) have emerged and gained significant attention. These models have, however, been shown to be notoriously difficult to train, which we hypothesize is partially due to the architectures, called ansatzes, that are hardly studied at this point. Therefore, in this paper, we take a step back and analyze ansatzes. We initially consider their expressivity, i.e., the space of operations they are able to express, and show that the closeness to being a 2-design, the primarily used measure, fails at capturing this property. Hence, we look for alternative ways to characterize ansatzes, unrelated to expressivity, by considering the local neighborhood of the model space, in particular, analyzing model distinguishability upon small perturbation of parameters. We derive an upper bound on their distinguishability, showcasing that QNNs using the Hardware Efficient Ansatz with few parameters are hardly discriminable upon update. Our numerical experiments support our bounds and further indicate that there is a significant degree of variability, which stresses the need for warm-starting or clever initialization. Altogether, our work provides an ansatz-centric perspective on training dynamics and difficulties in QNNs, ultimately suggesting that iterative training of small quantum models may not be effective, which contrasts their initial motivation.
@@ 1 Introduction
In light of these issues, we hypothesize that the architectures, called ansatzes, are part of the problem. QNN ansatzes consist of fixed and trainable gates, however, due to hardware limitations, currently mainly so-called Hardware Efficient Ansatzes (HEAs) (Kandala et al., 2017) are used, meaning they are built primarily based on hardware constraints. Therefore, we hypothesize that executing the circuit does not lead to a meaningful feature representation.
@@
In particular, in classical ML, a vanilla CNN is ultimately designed to consider local features in lower layers, before considering global features in deeper layers, independently of the dataset. Transformers (Vaswani et al., 2017) work similarly, however, the trainable one-qubit and fixed two-qubit gates that are used in QNNs do not allow drawing intuitive conclusions about how features are extracted.
Therefore, the goal of this paper is to analyze and characterize QNN architectures. We aim to take a first step at exploring ansatzes and the model space, initially by considering the expressivity of ansatzes. Using tools from Quantum Information Theory and extensive numerics, we later analyze how the quantum circuit, i.e., the applied operation, changes upon parameter update, thus, providing insights into the model space beyond ansatz expressivity. Our contributions are as follows.
  • Based on literature, we argue that closeness to a 2-design, a widely used measure for ansatz expressivity, does not adequately represent the possible set of unitary operations.
@@
  • We provide evidence that HEAs are hardly distinguishable upon parameter update in (1), random perturbations and (2), during training, even in early stages with larger updates.
  • While not strictly equal, the observed behavior and effects thereof have remarkable similarities to Barren Plateaus, insinuating two perspectives on similar trainability issues. The results suggest that the architectures, independently of data and loss, may be flawed and play a fundamental role in the trainability problems observed today.
The paper is structured as follows. In Section 2, we provide background information on QNNs, existing expressivity measures and related work [^1] We refer to Appendix A.1 for a brief overview of Quantum Computing.. In Section 3, we discuss the closeness to being a 2-design as a measure of expressivity and Section 4 introduces a measure of distinguishability of quantum channels. We elaborate on our experimental setup in Section 5, and present our results in Section 6. Section 7 discusses implications of our results and Section 8 concludes the paper.
@@ 2.1 Quantum Neural Networks
The small systems and the noise in today’s Noisy Intermediate-Scale Quantum (NISQ) technology (Preskill, 2018), prevents execution of large-scale circuits. Consequently, so-called Variational Quantum Circuits (VQC) are applied, which are hybrid models with classical and quantum parts.
Simply put, a QNN consists of (1) an input state |$\ket{\psi_{\bm{X}}}$, (2) an ansatz $U(\bm{\vartheta})$, (3) a loss function, and (4) an optimizer. The features are encoded onto the quantum state by means of a unitary feature encoding method $V(\bm{X})$, and the initial state is obtained as |$\ket{\psi_{\bm{X}}}=V(\bm{X})\ket{0}^{\otimes n}$, with $n$ as the number of qubits.
HEAs consists of $L$ layers of trainable one-qubit rotation gates $V_{i}(\vartheta_{i})$ and fixed two-qubit entangling gates $W_{i}$, with $i\in[1,L]$, expressed as Equation 1. A measurement is taken in the end, which is done through a Hamiltonian observable $\hat{O}$, and the result is expressed as the expectation value of the ansatz with respect to |$\ket{\psi_{\bm{X}}}$ and $\hat{O}$, as shown in Equation 2.
@@
2.2 Ansatz expressivity
We acknowledge a missing clear definition of expressivity in QML for now. It is sensible, therefore, to transfer it from classical ML, where model expressivity is the range of functions a model can compute (Raghu et al., 2017). Hence, it is a function of architecture $A$, input $x$ and parameters $\bm{\vartheta}$; $F_{A}(x;\bm{\vartheta})$. On the contrary, our work lies in data-agnostic characterization of models. To this end, we will only consider ansatz expressivity; the set of functions generated solely by an ansatz itself.
The most commonly used method to evaluate ansatz expressivity is to compare the probability distribution of the unitaries of the ansatz to the Haar measure [^2] We refer to Appendix A.3 for a short introduction to unitary groups and the Haar measure., the uniform distribution over all unitaries $\mathcal{U}(d)$ (Cerezo et al., 2021a). If the probability distribution of the ansatz follows the Haar measure up to the $t$-th moment, it is considered a t-design. Their construction requires exponential time (Harrow & Low, 2009), however, approximate t-designs can be built efficiently (Dankert et al., 2009).
@@ 2.3 Related work
BPs have been proven for deep (McClean et al., 2018) and shallow circuits with global measurement (Cerezo et al., 2021b), expressivity (Holmes et al., 2022; McClean et al., 2018), noise (Wang et al., 2021), entanglement (Ortiz Marrero et al., 2021) and it has been linked to cost concentration and narrow gorges (Arrasmith et al., 2022). Further, even shallow models without BPs have only a small fraction of local minima close to the global minimum (Anschuetz & Kiani, 2022) [^3] For further information, we refer to  Larocca et al. (2024), who published an excellent review on the topic..
The problem has been linked to the curse of dimensionality in Cerezo et al. (2024), therefore, limiting the accessible Hilbert space was proposed. However, they showed that such models can be simulated classically, given an initial data acquisition phase on a quantum computer. This was extended to a trade-off between trainability and dequantizability in Gil-Fuster et al. (2024).
Besides that, parameter initialization strategies have been proposed, with some even providing better bounds on gradient magnitudes (Rad et al., 2022; Wang et al., 2023; Kulshrestha & Safro, 2022; Grant et al., 2019), however, the results of Herbst et al. (2024) imply that it is not the type of statistical distribution, but rather the used parameter ranges that could lead to better starting points.
Further, tum advantage with respect to classical models. The closeness to being a 2-design is the most widely employed measure, however, others have been proposed. In particular, Equation 3 is related to the Frame Potential (Gross et al., 2007; Roberts & Yoshida, 2017; Nakaji & Yamamoto, 2021). Another approach is to analyze the entangling capability as in Sim et al. (2019), i.e., the entanglement of the produced states. Other measures are the Covering Number from  Du et al. (2022), analyzing Fourier coefficients (Schuld et al., 2021; Caro et al., 2021), or the Effective Dimension from Abbas et al. (2021). Further, one can consider unitary t-designs, i.e., consider the operations without the input state [^4] Unitary designs and the difference between unitary and state t-designs are discussed in Appendix A.3.. We employ the closeness to being a 2-design in our study, as it is frequently used, due to its intuitive interpretation. Beyond that, due to computational complexity, there is little research on properties of the architectures for now.
@@
Changes in quantum channel or loss landscape upon perturbing parameters has not yet been studied for QNNs. However, perturbation analysis and sensitivity analysis have emerged as tools in classical ML. Examples are perturbing the input for explainability (Ivanovs et al., 2021; Pizarroso et al., 2022; Fel et al., 2023), or checking for stability (Testa et al., 2024). Similarly, weights can be perturbed for trainability (Wen et al., 2018), or to increase robustness and generalization (Wu et al., 2020; He et al., 2019; Dumford & Scheirer, 2020). To the best of our knowledge, there is no research on perturbing parameters to check how the underlying function changes, as, due to the stochasticity of quantum computing, this research question seems to be much more relevant in the quantum realm.
@@
While 2-designs have turned out to be sufficient for many quantum information processing protocols (Holmes et al., 2022; Harrow & Low, 2009), in this section, we summarize and connect results from the literature to argue that closeness to a 2-design is an inadequate measure for model expressivity (i.e., the ability to sufficiently approximate the unitary group to a high degree).
@@
[…]
State t-design Welch bound: The choice of the polynomial function for state $t$-designs is presented in Appendix A.4. We state the following inequality for a state in a d-dimensional Hilbert space, |$\ket{\bm{\psi}_{i}}\in\mathbb{C}^{d}$ (cf. Appendix B.2 for the derivation):
$\sum_{1\leq j,k\leq n}|\langle\bm{\psi}_{j}|\bm{\psi}_{k}\rangle|^{2t}\geq\frac{n^{2}}{\binom{d+t-1}{t}}=\frac{n^{2}}{\frac{(d+t-1)(d+t-2).....d}{t!}}\ \ \forall t\in\mathbb{N}.$ (4)
It is clear that for an increasing $t$, the term in the denominator $\frac{(d+t-1)(d+t-2).....d}{t!}$ increases rapidly (as compared to Equation 24 for 2-designs), yielding a vanishing overlap between the state vector summands. This makes the inequality in Equation 4 more stringent for increasing $t$ values. This indicates that higher-order designs ($t>2$) approximating the Haar measure up to the $t$-th order impose stronger constraints, requiring greater separation between a sequence of state vectors $\{\bm{\psi}_{1},\bm{\psi}_{2},....,\bm{\psi}_{d}\}$ to maintain equality in the bounds. This means that 2-design models have far fewer degrees of freedom in building complex quantum circuits. On the contrary, t-designs ($t>2$) can more efficiently construct higher-order representations capturing more complex interactions between quantum states. Thus, evaluating Welch bounds for $2$-designs corroborates its inadequacy for constituting an expressive architecture.
An ansatz forming a 2-design is not expressive as following the Haar measure only up to the second moment constrains how spread-out (statistical spread) the unitaries must be, however, does not allow drawing conclusions about how densely it covers $\mathcal{U}(N)$. If one were to define expressivity as the capability of a model to represent all possible set of unitaries, it would be required to go way beyond second moments (for, e.g., Kurtosis, off-diagonal correlations, etc.) to ensure the necessary weights are captured where there is sufficiently dense population. As the moment operator requires tensor products proportional to the moments, this becomes computationally infeasible very quickly. It is not clear how to overcome the curse of dimensionality while proposing a measure that captures expressivity adequately; in fact, most tools from QIT suffer from this very phenomenon, speaking to the difficulty of designing such a metric. Further, there already exists a line of non-unitary VQA research (e.g., Cong et al. (2019); Deshpande et al. (2024)), that needs to be considered as well.
As expressivity is currently the most common metric used to describe the ansatzes, our work employs an alternative approach to analyze the model space, unrelated to expressivity, but highly relevant for trainability and training dynamics. Therefore, we focus on local neighborhoods of the model space, in particular, we study how the QNN changes upon perturbation of parameters. This allows analyzing training dynamics and tracking the extent to which the QNN changes during training.
@@ 4 Channel sensitivity
The diamond norm of a superoperator $T$ is defined as (Harrow & Low, 2009; Kitaev et al., 2002).
@@
Operationally, the diamond norm measures the maximum distinguishability between the output states of the two maps under any input state, i.e., for any input state we will be less or equally able to distinguish the output. A value of 00, means that they are indistinguishable, whereas a value of $2$ means they are perfectly distinguishable. Computing the diamond norm is non-trivial, however, can be reduced to a semi-definite program (Watrous, 2009; 2013).
We apply the diamond norm to evaluate the distinguishability of a parametrized unitary $U(\bm{\vartheta})$, representing a quantum channel, from the unitary $U(\bm{\vartheta}+\bm{\delta})$, where $\bm{\delta}$ is a perturbation of $\bm{\vartheta}$, as follows.
@@
We can provide an upper bound on the channel sensitivity through Taylor expansion, which is shown in Equation 7. Therefore, we establish a direct dependence of the distinguishability of the channels to the sum of changes applied [^6] For the precise mathematical derivation, we refer to Appendix C.. The bound holds if the Hermitian generators of the trainable gates are unitary as well, which is relevant for HEAs, as the trainable gates are exponentiations of the $X$, $Y$, and $Z$ Pauli gates. The precision of the bound depends on the magnitude of $\bm{\delta}$, however, considering that the models are trained iteratively in small steps, we expect the bound to hold.
$\begin{split}\|U(\bm{\vartheta})-U(\bm{\vartheta}+\bm{\delta})\|_{\Diamond}&=\|U(\bm{\vartheta})-(U(\bm{\vartheta})+\sum_{j=1}^{dim(\bm{\delta})}\delta_{j}\frac{\partial U(\bm{\vartheta})}{\partial\vartheta_{j}}+O(\bm{\delta}^{2}))\|_{\Diamond}\\ &\approx\|-\sum_{j=1}^{dim(\bm{\delta})}\delta_{j}\frac{\partial U(\bm{\vartheta})}{\partial\vartheta_{j}}\|_{\Diamond}\leq\frac{\sum_{j=1}^{dim(\bm{\delta})}|\delta_{j}|}{2}\end{split}$ (7)
Our bound shows that (assuming small parameters updates), there is a direct relationship between the number of parameters and the distinguishability of the operation. In particular, for small models, which are used extensively at the moment, it can be expected that subsequent unitaries during training are hardly distinguishable, hindering effective training. This is, to the best of our knowledge, the first result attempting to establish a connection between ansatz and trainability issues. To strengthen our results, we want to, in the following, support our bounds through numerical experiments.